OpenStudy (anonymous):

help! x-4/3 x 3/48x^2

OpenStudy (anonymous):

@kc_kennylau mind helping out?(:

OpenStudy (kc_kennylau):

can you rephrase the question using brackets

OpenStudy (anonymous):

yea one minute

OpenStudy (anonymous):

\[\frac{ x-4 }{ 3 } \times \frac{ 3 }{ 48x ^{2} }\]

OpenStudy (kc_kennylau):

well, the first step is to cancel the 3 :)

OpenStudy (anonymous):

yes I got it down to x-4/1 x 1/48x^2

OpenStudy (anonymous):

im stuck from there

OpenStudy (kc_kennylau):

\[\Large\frac{x-4}3\times\frac3{48x^2}=\frac{x-4}{48x^2}\]

OpenStudy (kc_kennylau):

This is the first step :)

OpenStudy (anonymous):

so now we simplify? and you smile a lot lol

OpenStudy (kc_kennylau):

And there is no second step :) I smile a lot because I am happy :D

OpenStudy (anonymous):

haha I see(: